# Mitigating the Accuracy-Robustness Trade-off via Multi-Teacher Adversarial Distillation

### v1

@@ 4 Experiments
In this section, we initially describe the experimental setting and then evaluate the Weighted Robust Accuracy of our MTARD and several baseline methods under prevailing white-box attacks and black-box attacks including transfer-based and square-based attacks. We also conduct a series of ablation studies to demonstrate the effectiveness of our method. Our code is available at https://github.com/zhaoshiji123/MTARD-extension.
@@ 4.1 Experimental Settings
We conduct our experiments on three datasets including CIFAR-10 [17], CIFAR-100, and Tiny-ImageNet [19], and consider the standard training method and several state-of-the-art methods as comparison methods: Adversarial Training method: SAT [25], Adversarial Robustness Distillation methods: ARD [9] and RSLAD [49]; Mitigating Trade-off methods: TRADES [45] and SCORE [26].
TABLE I: Performance of the teacher networks used in our experiments.
Dataset
Model
Clean Acc
FGSM
[…]
${}_{\scriptscriptstyle\rm{\infty}}$
AA
Type
CIFAR-10
ResNet-56
93.18%
19.23%
0%
0%
0%
0%
Clean
WideResNet-34-10
84.91%
61.14% 55.30% 56.61% 53.84% 52.53%
Adv
CIFAR-100
WideResNet-22-6
76.65%
4.85%
0%
0%
0%
0%
Clean
WideResNet-70-16
60.96%
35.89% 33.58% 33.99% 31.05% 30.03%
Adv
Tiny-ImageNet
PreActResNet-34
66.05%
2.50%
0.04%
0.09%
0%
0%
Clean
PreActResNet-34
52.76%
27.05% 24.00% 24.63% 20.07% 18.92%
Adv
@@
Student and Teacher Networks. Here we consider two student networks for CIFAR-10 and CIFAR-100 including ResNet-18 [13] and MobileNet-v2 [28]. For Tiny-ImageNet, we select PreActResNet-18 [14] and MobileNet-v2 following previous work. As for teacher models, we choose clean teacher networks including ResNet-56 for CIFAR-10, WideResNet-22-6 [44] for CIFAR-100, and PreActResNet-34 [14] for Tiny-ImageNet. Three robust teacher networks include WideResNet-34-10 for CIFAR-10, WideResNet-70-16 [11] for CIFAR-100, and PreActResNet-34 for Tiny-ImageNet. For CIFAR-10, WideResNet-34-10 is trained using TRADES [45]; For CIFAR-100, we use the WideResNet-70-16 model provided by Gowal et al. [11]; In Addition, for Tiny-ImageNet, we use PreActResNet-34 trained by TRADES [45] as the robust teacher. The whole teachers are pre-trained and will not be changed during the training process. The performance of these teacher models is shown in Table I.
TABLE II: White-box robustness on CIFAR-10, CIFAR-100, and Tiny-ImageNet datasets. The results of CIFAR-10 and CIFAR-100 are based on the ResNet-18, and the results of Tiny-ImageNet are based on the PreActResNet-18. All results are the best checkpoint based on W-Robust Acc.
CIFAR-10 CIFAR-100 Tiny-ImageNet
Attack
Defense
Clean
Robust
W-Robust
Clean
Robust
W-Robust
Clean
Robust
W-Robust
FGSM[10]
Natural
94.57%
18.60%
56.59%
75.18%
7.96%
41.57%
65.03%
1.73%
33.38%
SAT[25]
84.20%
55.59%
69.90%
56.16%
25.88%
41.02%
50.08%
25.35%
37.72%
TRADES[45]
83.00%
58.35%
70.68%
57.75%
31.36%
44.56%
48.45%
23.59%
36.02%
ARD[9]
84.11%
58.40%
71.26%
60.11%
33.61%
46.86%
53.22%
27.97%
40.60%
RSLAD[49]
83.99%
60.41%
72.20%
58.25%
34.73%
46.49%
48.78%
27.26%
38.02%
SCORE[26]
84.43%
59.84%
72.14%
56.40%
32.94%
44.67%
10.05%
7.80%
8.93%
MTARD(ours)
88.20%
61.42%
74.81%
65.08%
34.21%
49.65%
56.81%
28.12%
42.47%
[…]
Natural
94.57%
0%
47.29%
75.18%
0%
37.59%
65.03%
0.03%
32.53%
SAT[25]
84.20%
45.95%
65.08%
56.16%
21.18%
38.67%
50.08%
22.24%
36.16%
TRADES[45]
83.00%
52.35%
67.68%
57.75%
28.05%
42.90%
48.45%
21.59%
35.02%
ARD[9]
84.11%
50.93%
67.52%
60.11%
29.4%%
44.76%
53.22%
24.92%
39.07%
RSLAD[49]
83.99%
53.94%
68.97%
58.25%
31.19%
44.72%
48.78%
25.00%
36.89%
SCORE[26]
84.43%
53.72%
69.08%
56.40%
30.27%
43.34%
10.05%
7.65%
8.85%
MTARD(ours)
88.20%
51.68%
69.94%
65.08%
28.50%
46.79%
56.81%
23.93%
40.37%
[…]
Natural
94.57%
0%
47.29%
75.18%
0%
37.59%
65.03%
0.05%
32.54%
SAT[25]
84.20%
48.12%
66.16%
56.16%
22.02%
39.09%
50.08%
23.05%
36.57%
TRADES[45]
83.00%
53.83%
68.42%
57.75%
28.88%
43.32%
48.45%
22.09%
35.27%
ARD[9]
84.11%
52.96%
68.54%
60.11%
30.51%
45.31%
53.22%
25.71%
39.47%
RSLAD[49]
83.99%
55.73%
69.86%
58.25%
32.05%
45.15%
48.78%
25.45%
37.12%
SCORE[26]
84.43%
55.21%
69.82%
56.40%
30.56%
43.48%
10.05%
7.67%
8.86%
MTARD(ours)
88.20%
54.40%
71.30%
65.08%
29.94%
47.51%
56.81%
24.94%
40.88%
${}_{\scriptscriptstyle\rm{\infty}}$[3]
Natural
94.57%
0%
47.29%
75.18%
0%
37.59%
65.03%
0%
32.52%
SAT[25]
84.2%
45.97%
65.09%
56.16%
20.90%
38.53%
50.08%
20.48%
35.28%
TRADES[45]
83.00%
50.23%
66.62%
57.75%
24.19%
40.97%
48.45%
17.33%
32.89%
ARD[9]
84.11%
50.15%
67.13%
60.11%
27.56%
43.84%
53.22%
21.41%
37.32%
RSLAD[49]
83.99%
52.67%
68.33%
58.25%
28.21%
43.23%
48.78%
20.87%
34.83%
SCORE[26]
84.43%
50.46%
67.45%
56.40%
26.30%
41.35%
10.05%
6.19%
8.13%
MTARD(ours)
88.20%
49.88%
69.04%
65.08%
25.45%
45.27%
56.81%
19.69%
38.25%
AutoAttack[6]
Natural
94.57%
0%
47.29%
75.18%
0%
37.59%
65.03%
0%
32.52%
SAT[25]
84.20%
43.49%
63.85%
56.16%
19.76%
37.96%
50.08%
18.54%
34.31%
TRADES[45]
83.00%
49.03%
66.02%
57.75%
23.26%
40.51%
48.45%
16.66%
32.56%
ARD[9]
84.11%
48.20%
66.16%
60.11%
25.71%
42.91%
53.22%
19.91%
36.57%
RSLAD[49]
83.99%
50.98%
67.49%
58.25%
26.50%
42.38%
48.78%
17.11%
32.95%
SCORE[26]
84.43%
49.25%
66.84%
56.40%
24.74%
40.57%
10.05%
5.97%
8.01%
MTARD(ours)
88.20%
47.44%
67.82%
65.08%
23.98%
44.53%
56.81%
17.71%
37.26%
TABLE III: White-box robustness on CIFAR-10, CIFAR-100, and Tiny-ImageNet datasets. The result of three datasets is trained based on the MobileNet-v2. All results are the best checkpoint based on W-Robust Acc.
CIFAR-10 CIFAR-100 Tiny-ImageNet
Attack
Defense
Clean
Robust
W-Robust
Clean
Robust
W-Robust
Clean
Robust
W-Robust
FGSM[10]
Natural
93.35%
12.22%
52.79%
74.86%
5.94%
40.40%
62.09%
0.72%
31.41%
SAT[25]
83.87%
55.89%
69.88%
59.19%
30.88%
45.04%
49.03%
23.38%
36.21%
TRADES[45]
77.95%
53.75%
65.85%
55.41%
30.28%
42.85%
43.81%
20.10%
31.96%
ARD[9]
83.43%
57.03%
70.23%
60.45%
32.77%
46.61%
45.53%
22.88%
33.21%
RSLAD[49]
83.20%
59.47%
71.34%
59.01%
33.88%
46.45%
45.69%
24.09%
34.89%
SCORE[26]
82.32%
58.43%
70.38%
49.38%
29.28%
39.33%
28.27%
17.47%
22.87%
MTARD(ours)
88.70%
58.78%
73.74%
66.13%
34.36%
50.25%
52.98%
25.60%
39.29%
[…]
Natural
93.35%
0%
46.68%
74.86%
0%
37.43%
62.09%
0%
31.05%
SAT[25]
83.87%
46.84%
65.36%
59.19%
25.64%
42.42%
49.03%
20.31%
34.67%
TRADES[45]
77.95%
49.06%
63.51%
55.41%
23.33%
39.37%
43.81%
18.16%
30.99%
ARD[9]
83.43%
49.50%
66.47%
60.45%
28.69%
44.57%
45.53%
20.43%
32.98%
RSLAD[49]
83.20%
53.25%
68.23%
59.01%
30.19%
44.60%
45.69%
22.30%
34.00%
SCORE[26]
82.32%
53.42%
67.87%
49.38%
27.03%
38.21%
28.27%
16.48%
22.38%
MTARD(ours)
88.70%
48.50%
68.60%
66.13%
28.47%
47.30%
52.98%
21.58%
37.28%
[…]
Natural
93.35%
0%
46.68%
74.86%
0%
37.43%
62.09%
0%
31.05%
SAT[25]
83.87%
49.14%
66.51%
59.19%
26.96%
43.08%
49.03%
21.15%
35.09%
TRADES[45]
77.95%
50.27%
64.11%
55.41%
28.42%
41.92%
43.81%
18.36%
31.09%
ARD[9]
83.43%
51.70%
67.57%
60.45%
29.63%
45.04%
45.53%
21.00%
33.27%
RSLAD[49]
83.20%
54.76%
68.98%
59.01%
31.19%
45.10%
45.69%
22.74%
34.22%
SCORE[26]
82.32%
54.46%
68.39%
49.38%
27.53%
38.46%
28.27%
16.69%
22.48%
MTARD(ours)
88.70%
51.11%
69.91%
66.13%
29.82%
47.98%
52.98%
22.58%
37.78%
${}_{\scriptscriptstyle\rm{\infty}}$[3]
Natural
93.35%
0%
46.68%
74.86%
0%
37.43%
62.09%
0%
31.05%
SAT[25]
83.87%
46.62%
65.25%
59.19%
25.01%
42.10%
49.03%
18.69%
33.86%
TRADES[45]
77.95%
46.06%
62.01%
55.41%
27.72%
41.57%
43.81%
13.47%
28.66%
ARD[9]
83.43%
48.96%
66.20%
60.45%
26.55%
43.50%
45.53%
16.81%
31.17%
RSLAD[49]
83.20%
51.78%
67.49%
59.01%
27.98%
43.50%
45.69%
18.63%
32.16%
SCORE[26]
82.32%
49.18%
65.75%
49.38%
23.29%
36.34%
28.27%
13.25%
20.76%
MTARD(ours)
88.70%
47.09%
67.90%
66.13%
26.50%
46.32%
52.98%
18.08%
35.53%
AutoAttack[6]
Natural
93.35%
0%
46.68%
74.86%
0%
37.43%
62.09%
0%
31.05%
SAT[25]
83.87%
43.43%
63.65%
59.19%
22.84%
41.02%
49.03%
17.35%
33.19%
TRADES[45]
77.95%
45.18%
61.66%
55.41%
22.61%
39.01%
43.81%
12.82%
28.32%
ARD[9]
83.43%
46.60%
65.02%
60.45%
24.86%
42.66%
45.53%
15.40%
30.47%
RSLAD[49]
83.20%
50.23%
66.72%
59.01%
26.33%
42.67%
45.69%
17.09%
31.39%
SCORE[26]
82.32%
48.39%
65.36%
49.38%
21.92%
35.65%
28.27%
12.30%
20.29%
MTARD(ours)
88.70%
44.93%
66.82%
66.13%
24.01%
45.07%
52.98%
15.68%
34.33%
Training Setting. We train the student using Stochastic Gradient Descent (SGD) optimizer with an initial learning rate of 0.1, a momentum of 0.9, and a weight decay of 2e-4. For MTARD, the weight loss learning rate $r_{w}$ is initially set as 0.025; meanwhile, the $\tau_{nat}$ and $\tau_{adv}$ are initially set as 1, and the temperature learning rate $r_{\tau}$ is initially set as 0.001. the student’s temperature $\tau_{s}$ is set to constant 1. The maximum and minimum values of temperature are 10 and 1 respectively. For CIFAR-10 and CIFAR-100, we set the total number of training epochs to 300. The learning rate is divided by 10 at the 215-th, 260-th, and 285-th epochs; For Tiny-ImageNet, we set the total number of training epochs to 100, and the learning rate is divided by 10 at the 75-th and 90-th epochs. We set the batch size to 128 (CIFAR-10 and CIFAR-100) and 32 (Tiny-ImageNet), and $\beta$ is set to 1. For the inner maximization of MTARD, we use a 10-step PGD with a random start size of 0.001 and a step size of 2/255. All training perturbation in the maximization is bounded to the $L_{\infty}$ norm $\epsilon$ = 8/255. For the training of the natural model, we train the networks for 100 epochs, and the learning rate is divided by 10 at the 75-th and 90-th epochs.
For the comparison methods, we follow SAT, TRADES, SCORE, ARD, and RSLAD original settings if without additional instruction. The SCORE version is the TRADES+LSE without additional data. For ARD, we use the same robust teacher as RSLAD and MTARD. The temperature $\tau$ of ARD is set to 30 on CIFAR-10 while set to 5 on CIFAR-100 and Tiny-ImageNet, and the $\alpha$ is set to 0.95 following [9] on CIFAR-100 and Tiny-ImageNet. The learning rate of SAT and SCORE for Tiny-ImageNet is 0.01 to achieve better performance.
[…]
Figure 3: Ablation study with ResNet-18 student trained using variants of our MTARD and Baseline method on CIFAR-10. NLB and EBB represent Normalization Loss Balance (NLB) and Entropy-Based Balance (EBB). MTARD is our final version, which represents Baseline+NLB+EBB. All the results are the best checkpoint based on W-Robust Acc.

### v3

@@ 4 Experiments
In this section, we initially describe the experimental setting and then evaluate the Weighted Robust Accuracy of our B-MTARD and several baseline methods under prevailing white-box attacks and black-box attacks including transfer-based and query-based attacks. We also conduct a series of ablation studies to demonstrate the effectiveness.
@@ 4.1 Experimental Settings
We conduct experiments on three datasets including CIFAR-10 [22], CIFAR-100, and Tiny-ImageNet [24]. We apply the standard training method and several state-of-the-art methods for comparison: Adversarial Training method: SAT [31], Adversarial Robustness Distillation methods: ARD [13], RSLAD [61], Fair-ARD [49], and ABSLD [58]; Mitigating Trade-off methods: TRADES [55] and SCORE [32]. Meanwhile, we also show the results of our conference version for comparison: MTARD [59].
@@
Student and Teacher Networks. Here we consider two student networks for CIFAR-10 and CIFAR-100 including ResNet-18 [17] and MobileNet-v2 [34]. For Tiny-ImageNet, we select PreActResNet-18 [18] and MobileNet-v2 following previous work. As for teacher models, we choose clean teacher networks including ResNet-56 for CIFAR-10, WideResNet-22-6 [54] for CIFAR-100, and PreActResNet-34 [18] for Tiny-ImageNet. Three robust teacher networks include WideResNet-34-10 for CIFAR-10, WideResNet-70-16 [15] for CIFAR-100, and PreActResNet-34 for Tiny-ImageNet. For CIFAR-10, WideResNet-34-10 is trained using TRADES [55]; For CIFAR-100, we use the WideResNet-70-16 model provided by Gowal et al. [15]; In Addition, for Tiny-ImageNet, we use PreActResNet-34 trained by TRADES [55] as the robust teacher. The teachers are pre-trained and will not be changed during the training process. The performance of these teacher models is shown in Table II.
Training Setting. We train the student using Stochastic Gradient Descent (SGD) optimizer with an initial learning rate of 0.1, a momentum of 0.9, and a weight decay of 2e-4. For B-MTARD, the weight loss learning rate $r_{w}$ is initially set as 0.025; the temperature learning rate $r_{\tau}$ is initially set as 0.001. the student’s temperature $\tau_{s}$ is set to constant 1. Meanwhile, the $\tau_{nat}$ and $\tau_{adv}$ are initially set as 1 without additonal instruction, and the $\tau_{nat}$ and $\tau_{adv}$ are initially set as 2 for MobileNet-v2 on CIFAR-10. The maximum and minimum values of temperature are 10 and 1, respectively. For CIFAR-10 and CIFAR-100, we set the total number of training epochs to 300. The learning rate is divided by 10 at the 215-th, 260-th, and 285-th epochs; For Tiny-ImageNet, we set the total number of training epochs to 100, and the learning rate is divided by 10 at the 75-th and 90-th epochs. We set the batch size to 128 (CIFAR-10 and CIFAR-100) and 32 (Tiny-ImageNet), and $\beta$ is set to 1. For the inner maximization of B-MTARD, we use a 10-step PGD with a random start size of 0.001 and a step size of 2/255. All training perturbation in the maximization is bounded to the $L_{\infty}$ norm $\epsilon$ = 8/255. For the training of the natural model, we train the networks for 100 epochs, and the learning rate is divided by 10 at the 75-th and 90-th epochs.
For the comparison methods, we follow SAT, TRADES, SCORE, ARD, RSLAD, Fair-ARD, and ABSLD original settings without additional instruction. The SCORE version is the TRADES+LSE without additional data. For ARD, we use the same robust teacher as RSLAD and MTARD. The temperature $\tau$ of ARD is set to 30 on CIFAR-10 while set to 5 on CIFAR-100 and Tiny-ImageNet, and the $\alpha$ is set to 0.95 following [13] on CIFAR-100 and Tiny-ImageNet. The learning rate of SAT and SCORE for Tiny-ImageNet is 0.01 to achieve better performance.
Evaluation Setting. The same as previous studies, we evaluate the models against white box adversarial attacks: FGSM [14], PGDsat[31], PGDtrades[55], CW[4], which are commonly used adversarial attacks in adversarial robustness evaluation. Meanwhile, we also apply a strong attack: AutoAttack (AA) [10] (an evaluation tool in RobustBench [8]) to evaluate the robustness. The step sizes of PGDsat and PGDtrades are 2/255 and 0.003, respectively, and the step is 20. The total step of CW is 30. Maximum perturbation is bounded to the $L_{\infty}$ norm $\epsilon$ = 8/255 for all attacks. Meanwhile, we conduct a black-box evaluation, which includes the transfer-based attack and query-based attack to test the robustness in a near-real environment. As for the transfer-based attack, we choose robust teachers (WideResNet-34-10 for CIFAR-10, WideResNet-70-16 for CIFAR-100, and PreActResNet-34 for Tiny-ImageNet) as the surrogate models to produce adversarial example against the PGDtrades and CW attack; As for the query-based attack, we choose a score-based attack: Square Attack [2] and a decision-based attack: RayS Attack [6], and the queries are 100.
Figure 3: Ablation study with ResNet-18 student trained using variants of our B-MTARD and Baseline method on CIFAR-10. NLB and EBB represent Normalization Loss Balance (NLB) and Entropy-Based Balance (EBB). B-MTARD is our final version, which represents Baseline+NLB+EBB. The Baseline+NLB is the result in our ECCV version [59]. All the results are the best checkpoint based on W-Robust Acc.
@@
PGDtrades[55] CW[4] Square Attack[2] RayS Attack[6]
DataSet Defense Clean Robust W-Robust Robust W-Robust Robust W-Robust Robust W-Robust
CIFAR-10 SAT[31] 84.20% 64.84% 74.52% 63.84% 74.02% 72.48% 78.34% 74.13% 79.17%
TRADES[55] 83.00% 63.61% 73.31% 62.83% 72.92% 72.49% 77.75% 73.59% 78.30%
ARD[13] 84.11% 63.50% 73.81% 62.86% 73.49% 74.60% 79.36% 75.92% 80.02%
RSLAD[61] 83.99% 63.96% 73.98% 63.05% 73.52% 72.47% 78.23% 75.60% 79.80%
SCORE[32] 84.43% 64.34% 74.39% 63.79% 74.11% 74.68% 79.56% 75.87% 80.15%
Fair-ARD[49] 83.41% 63.31% 73.36% 62.76% 73.09% 74.08% 78.75% 75.38% 79.40%
ABSLD[58] 83.21% 62.91% 73.06% 62.19% 72.70% 73.93% 78.57% 74.90% 79.06%
MTARD[59] 87.36% 65.26% 76.31% 64.58% 75.97% 78.58% 82.97% 80.13% 83.75%
B-MTARD (ours) 88.20% 65.29% 76.75% 64.64% 76.42% 79.82% 84.01% 80.98% 84.59%
CIFAR-100 SAT[31] 56.16% 38.10% 47.13% 39.42% 47.79% 40.05% 48.11% 40.92% 48.54%
TRADES[55] 57.75% 38.20% 47.98% 38.63% 48.19% 42.51% 50.13% 42.77% 50.26%
ARD[13] 60.11% 39.53% 49.82% 38.85% 49.48% 47.20% 53.66% 48.04% 56.08%
RSLAD[61] 58.25% 39.93% 49.09% 39.67% 48.96% 45.32% 51.79% 45.77% 52.01%
SCORE[32] 56.40% 39.09% 47.75% 40.02% 48.21% 43.98% 50.34% 44.57% 50.49%
Fair-ARD[49] 57.81% 39.56% 48.69% 39.10% 48.46% 44.99% 51.40% 45.54% 51.68%
ABSLD[58] 56.77% 38.42% 47.60% 38.11% 47.44% 44.09% 50.43% 44.57% 50.67%
MTARD[59] 64.30% 41.46% 52.88% 41.18% 52.74% 48.13% 56.22% 49.73% 57.02%
B-MTARD (ours) 65.08% 42.11% 53.60% 41.35% 53.22% 49.40% 57.24% 50.66% 57.87%
Tiny-ImageNet SAT[31] 50.08% 33.40% 41.74% 33.20% 41.63% 38.72% 44.40% 39.35% 44.72%
TRADES[55] 48.45% 31.01% 39.73% 30.72% 39.59% 36.58% 42.52% 37.22% 42.84%
ARD[13] 53.22% 34.74% 43.98% 33.32% 43.27% 42.58% 47.90% 42.87% 48.05%
RSLAD[61] 48.78% 32.85% 40.82% 32.09% 40.44% 37.64% 43.21% 38.15% 43.47%
SCORE[32] 10.05% 8.74% 9.40% 8.82% 9.44% 8.67% 9.36% 8.70% 9.38%
Fair-ARD[49] 46.64% 31.58% 39.11% 31.38% 39.01% 35.81% 41.23% 36.44% 41.54%
ABSLD[58] 47.21% 31.84% 39.53% 31.66% 39.44% 36.77% 41.99% 37.42% 42.32%
MTARD[59] 52.98% 34.48% 43.73% 33.80% 43.39% 41.70% 47.34% 42.55% 47.77%
B-MTARD (ours) 56.81% 36.65% 46.73% 33.40% 45.11% 44.46% 50.64% 45.51% 51.16%
TABLE VI: Black-box robustness on CIFAR-10, CIFAR-100, and Tiny-ImageNet datasets. The results of three datasets are trained based on the MobileNet-v2. All results are the best checkpoint based on W-Robust Acc.
PGDtrades[55] CW[4] Square Attack[2] RayS Attack[6]
DataSet Defense Clean Robust W-Robust Robust W-Robust Robust W-Robust Robust W-Robust
CIFAR-10 SAT[31] 83.87% 64.66% 74.27% 64.24% 74.06% 73.01% 78.44% 74.72% 79.30%
TRADES[55] 77.95% 61.04% 69.50% 60.66% 69.31% 67.43% 72.69% 68.91% 73.43%
ARD[13] 83.43% 63.28% 73.36% 62.83% 73.13% 73.26% 78.35% 74.62% 79.03%
RSLAD[61] 83.20% 64.33% 73.77% 63.45% 73.33% 73.07% 78.14% 74.49% 78.85%
SCORE[32] 82.32% 63.84% 73.08% 63.03% 72.68% 72.41% 77.37% 73.45% 77.89%
Fair-ARD[49] 82.65% 62.62% 72.64% 62.00% 72.32% 72.44% 77.55% 73.50% 78.08%
ABSLD[58] 82.50% 62.91% 72.71% 61.87% 72.19% 72.05% 77.28% 73.29% 77.90%
MTARD[59] 89.26% 66.30% 77.78% 65.68% 77.47% 79.13% 84.20% 80.91% 85.09%
B-MTARD (ours) 89.09% 66.47% 77.78% 65.96% 77.53% 79.61% 84.35% 81.14% 85.12%
CIFAR-100 SAT[31] 59.19% 40.70% 49.95% 40.97% 50.08% 44.51% 51.85% 45.59% 52.39%
TRADES[55] 55.41% 37.76% 46.59% 38.02% 46.72% 40.71% 48.06% 41.23% 48.32%
ARD[13] 60.45% 39.15% 49.80% 38.53% 49.49% 46.95% 53.70% 47.91% 54.18%
RSLAD[61] 59.01% 40.32% 49.67% 39.92% 49.47% 45.66% 52.34% 46.33% 52.67%
SCORE[32] 49.38% 40.00% 44.69% 36.91% 43.15% 37.94% 43.66% 38.04% 43.71%
Fair-ARD[49] 59.18% 40.55% 49.87% 40.14% 49.66% 45.29% 52.24% 46.00% 52.59%
ABSLD[58] 56.67% 38.08% 47.38% 38.18% 47.43% 43.23% 49.95% 43.60% 50.14%
MTARD[59] 67.01% 43.23% 55.12% 42.92% 54.97% 50.64% 58.83% 52.49% 59.75%
B-MTARD (ours) 66.13% 42.67% 54.40% 42.04% 54.09% 50.83% 58.48% 51.84% 58.99%
Tiny-ImageNet SAT[31] 49.03% 33.47% 41.25% 33.13% 41.08% 37.95% 43.49% 38.32% 43.68%
TRADES[55] 43.81% 28.35% 36.08% 28.64% 36.23% 32.39% 38.10% 32.37% 38.09%
ARD[13] 45.53% 30.73% 38.13% 30.23% 37.88% 34.60% 40.07% 35.13% 40.33%
RSLAD[61] 45.69% 31.20% 38.45% 31.10% 38.40% 35.18% 40.44% 35.57% 40.63%
SCORE[32] 28.27% 21.82% 25.05% 22.19% 25.23% 22.16% 25.22% 22.41% 25.34%
Fair-ARD[49] 47.24% 31.80% 39.52% 31.40% 39.32% 36.53% 41.89% 37.22% 42.23%
ABSLD[58] 48.08% 32.89% 40.49% 32.43% 40.26% 38.18% 43.13% 38.48% 43.28%
MTARD[59] 50.50% 32.75% 41.63% 32.05% 41.28% 38.88% 44.69% 40.01% 45.26%
B-MTARD (ours) 52.98% 34.25% 43.62% 33.50% 43.24% 40.62% 46.80% 41.44% 47.21%